# Edge-Spreading Raptor-Like LDPC Codes for 6G Wireless Systems

### v1

@@ Abstract
Next-generation channel coding has stringent demands on throughput, energy consumption, and error rate performance while maintaining key features of 5G New Radio (NR) standard codes such as rate compatibility, which is a significant challenge. Due to excellent capacity-achieving performance, spatially-coupled low-density parity-check (SC-LDPC) codes are considered a promising candidate for next-generation channel coding. In this paper, we propose an SC-LDPC code family called edge-spreading Raptor-like (ESRL) codes for 6G. Unlike other SC-LDPC codes that adopt the structure of existing rate-compatible LDPC block codes before coupling, ESRL codes maximize the possible locations of edge placement and focus on constructing an optimal coupled matrix. Moreover, a new graph representation called the unified graph is introduced. This graph offers a global perspective on ESRL codes and identifies the optimal edge reallocation to optimize the spreading strategy. We conduct comprehensive comparisons of ESRL codes and 5G-NR LDPC codes. Simulation results demonstrate that when all decoding parameters and complexity are the same, ESRL codes have obvious advantages in error rate performance and throughput compared to 5G-NR LDPC codes, making them a promising solution towards next-generation channel coding.
@@ I Introduction
Next-generation channel coding targets a peak data rate of $200$ Gbps to $1$ Tbps with an energy efficiency target of around $1$ pJ/bit [1], representing 10×10$\times$10 × to 100×100$\times$100 × more stringent requirements compared to 5G New Radio (NR) [2]. As standards continue to evolve, channel coding also needs to meet diverse key performance indicators (KPIs), including higher throughput, enhanced reliability, lower power consumption, reduced computational complexity, rate compatibility, and new capabilities [3], leading to a significant challenge.
@@
As a key technique in 5G-NR [4], low-density parity-check (LDPC) codes [5] have achieved significant success due to capacity-approaching performance [6] and parallel processing capabilities under belief propagation (BP) decoding [7]. For finite-length LDPC codes, though BP offers the advantage of reduced decoding complexity, the BP decoding threshold is sub-optimal compared to the maximum a-posteriori probability (MAP) threshold. Hence, spatially-coupled LDPC (SC-LDPC) codes (also called LDPC convolutional codes) were proposed in [8] and [9], by superimposing LDPC block codes with a convolutional structure. In general, block codes have a simple structure that is easy to implement in hardware, but they have increased complexity and latency when handling frames with large blocklengths. In contrast, convolutional codes are effective at decoding long frames but suffer from sequential processing of windowed decoding that introduces decoding latency. SC-LDPC codes, offering the best of both worlds, achieve the MAP threshold of the underlying block ensembles using BP decoding on a coupled ensemble [9, 10]. Besides, the compact and regular structure of underlying blocks also simplifies the decoding of long frames, and state-of-the-art SC-LDPC decoder implementations have reported high throughput in the range of hundreds of Gbps [11]. These ideas have attracted widespread interest for next-generation wireless and optical fiber communications [3, 12].
@@
For wireless channels, rate compatibility and blocklength scalability are necessary features for channel coding. To address the time-varying nature of wireless channels, rate-compatible channel codes support incremental redundancy transmission, thereby maximizing the throughput under hybrid automatic repeat-request (HARQ) schemes [13]. Specifically, a rate-compatible code family is a nested sequence of codes with different rates, where higher rate codes are embedded into lower rate codes [14]. An initial codeword is transmitted at a high rate, and in case of erroneous reception or failed decoding, additional partial redundancy is sent until all redundancy of the mother code is used up. Numerous works on rate-compatible LDPC codes have been proposed in [14, 15, 16, 17] using puncturing and extension. Moreover, the design of LDPC codes can begin with the construction of a relatively small bipartite graph called the protograph [18]. A well-designed protograph applies a copy-and-permute operation (also called lifting) to obtain larger graphs with nodes, thus leading to longer-blocklength LDPC codes [19]. In this context, a protograph-based Raptor-like (PBRL) LDPC code family that meets the above requirements is a very popular one [17]. Sharing many similarities with LDPC codes, the introduction of Luby transform (LT) codes [20] and Raptor codes [21], which can virtually generate any desired rate, form a novel coding paradigm of rateless codes. Subsequently, PBRL codes restrict the protograph to a Raptor-like structure, achieving both rate compatibility and blocklength adjustment simultaneously. Hence, PBRL LDPC codes have been adopted into 5G-NR successfully [4]. However, the complexity of BP decoding and hardware increases as the blocklength grows in 5G-NR. As indicated by the state-of-the-art 5G-NR LDPC decoder implementation [22], the memory overhead required to accommodate the maximum blocklength occupies a significant part of the decoder core area (over $60\%$) and results in complex routing. This limits peak throughput to around $20$ Gbps for 5G-NR LDPC codes, far away from stringent demands beyond 5G.
@@
Given the advantages of SC-LDPC codes over traditional LDPC block codes, developing a rate-compatible SC-LDPC code family has emerged as a promising solution towards next-generation channel coding. Note that SC-LDPC codes can be constructed from LDPC block codes using an edge-spreading technique [23]. Several SC-LDPC code constructions tailored to a specific rate were presented in [24, 25, 26, 27, 28, 29], where most of them focus on optimizing edge spreading starting from a given LDPC block code, mitigating the emergence of harmful graph structures. For instance, [26] proposed a heuristic search combined with a check-and-flip process to construct short-cycle-free SC-LDPC codes by generating multiple auxiliary matrices. However, the pseudo-random nature of [26] leads to significant complexity, particularly as the matrix size and coupling width increase. To address this issue, an edge-spreading matrix [30] has been employed in recent works [25, 27, 28, 29] to provide a more compact representation of edge spreading. This enables a simple method to identify harmful structures in SC-LDPC codes, regardless of the increase in coupling width. Sima et al. [25] imposed constraints on the edge-spreading matrix to reduce dominant trapping sets and avoid error floors, and [28] analyzed the relationship between short cycles and edge spreading to design short-cycle-free SC-LDPC codes with minimal coupling width. To further incorporate rate compatibility [31, 32, 33, 34, 35], Si et al. introduced the first rate-compatible SC-LDPC code to achieve the capacity of the binary erasure channel (BEC) [31]. Zhou et al. designed a family of robust rate-compatible punctured SC-LDPC codes by ensuring the recoverability of punctured bits and avoiding harmful graph structures. Protograph-based rate-compatible SC-LDPC codes were proposed in [33, 34, 35] following a simple idea. Namely, a conventional PBRL protograph is first designed, and then the edge-spreading technique is directly applied to construct a coupled code, with both procedures being independent. Notably, [35] achieved performance gains by simply splitting the 5G-NR LDPC matrix as the foundation for SC-LDPC codes.
@@
However, these SC-LDPC works either focus on optimizing code constructions tailored to a specific rate [24, 25, 26, 27, 28, 29] or they apply the structure of existing rate-compatible LDPC block codes to generate SC-LDPC codes without considering the influence of the coupling on performance [33, 34, 35], thus not exploring the full potential of SC-LDPC codes. Consequently, there is a need for a holistic design method to construct SC-LDPC codes that simultaneously meet diverse demands, including rate compatibility, protograph design with coupling, threshold analysis, removal of harmful graph structures, and high throughput potential. Moreover, it is necessary to conduct comprehensive comparisons of the designed SC-LDPC codes and standard LDPC codes, especially when all decoding parameters and complexity are consistent. This would provide a clear assessment of the potential of SC-LDPC codes for next-generation channel coding.
Contributions: This paper proposes an SC-LDPC code family called edge-spreading Raptor-like (ESRL) LDPC codes. Unlike PBRL codes [17], ESRL codes maximize the possible locations of edge placement before coupling and focus on constructing an optimal coupled matrix. Hence, ESRL codes show outstanding performance while maintaining extensive rate compatibility. The structure, properties, and code profile of ESRL codes are discussed in detail.
Moreover, we propose a novel graph representation called the unified graph. Unlike other edge-spreading optimizations involved with complex processing [27, 28, 29, 30], this graph innovatively offers a global perspective on ESRL codes, identifying the optimal edge reallocation to optimize the spreading strategy. We also introduce modified coupled reciprocal channel approximation (RCA) to analyze the decoding threshold of ESRL codes, demonstrating that ESRL codes enjoy better thresholds across a wide range of rates. Based on these tools, a complete design flow of ESRL codes is presented.
This paper also proposes a high-throughput decoding algorithm tailored to ESRL codes, called semi-layered multi-engine (SLME) decoding. Based on features of ESRL codes, SLME decoding decomposes the original BP decoding into multiple identical sub-decoders that operate in parallel in a more compact manner compared to [12]. This approach significantly enhances the decoding parallelism for ESRL codes.
Finally, we provide a design example of ESRL codes and conduct comprehensive comparisons with 5G-NR LDPC codes [4]. As summarized in Fig. 1, ESRL codes preserve several advantages and properties of 5G-NR LDPC codes while further enhancing error rate performance and throughput. When all decoding parameters and complexity are the same, at the rate of $0.52$, ESRL codes outperform 5G-NR LDPC codes by $0.95$ dB at five iterations or enhance peak throughput by $50.1\%$ using a similar hardware architecture [22].
@@
Outline: The remainder of this paper is organized as follows. Section II introduces the symbol definitions and reviews in compact form the preliminaries of LDPC codes. Section III introduces the structure, properties, and code profile of ESRL codes. Section IV outlines the design and optimization procedures for ESRL codes, including the unified graph and modified coupled RCA. In Section V, we present the puncturing scheme and SLME decoding for ESRL codes. Section VI provides a design example of ESRL codes and compares it with 5G-NR LDPC codes. Finally, Section VII concludes this paper.
@@ II Preliminaries
LDPC block codes are specified by a sparse $M\times N$ parity-check matrix (PCM) $\mathbf{H}$, which can be further described by a bipartite Tanner graph [36] with the information length of $K=N-M$. If $\mathbf{H}_{c,v}=1$ for $0\leq c<M$ and $0\leq v<N$, the $c$-th check node (CN) is connected to the $v$-th variable node (VN) on the Tanner graph. We use $\mathbb{V}_{c}$ to represent the set of adjacent VNs for the $c$-th CN and we use $\mathbb{C}_{v}$ to denote the set of neighbors for the $v$-th VN. Let $|\mathbb{V}_{c}|=d_{c}$ and $|\mathbb{C}_{v}|=d_{v}$ refer to the row degree of the $c$-th CN and the column degree of the $v$-th VN, respectively. Due to the quasi-cyclic (QC) property [19], a QC-LDPC code can be described by a more structured $m\times n$ protomatrix $\mathbf{B}$ with the information length of $k=n-m$. where each entry $\mathbf{B}_{c,v}$ can be extended by a $Z\times Z$ cyclic-shifting identity matrix, with $Z$ being the lifting size. To focus on the protomatrix design in this paper, $\mathbf{B}_{c,v}$ denotes only the number of connected edges between the $c$-th CN and the $v$-th VN instead of a specific cyclic shifting value. Similarly, $\mathbf{B}$ can be transformed into a protograph $\mathcal{G}$ [18], defined as $\mathcal{G}\triangleq\{\mathbb{V}\cup\mathbb{C},\mathbb{E}\}$, where $\mathbb{E}$ is the set of all edges in $\mathcal{G}$. Let $e_{c}^{v}$ denote the directed edge from the $c$-th CN to the $v$-th VN, and $e_{v}^{c}$ is the same edge with the opposite direction. Note that all graph concepts are based on protographs throughout this paper and we omit the term ‘QC’ for simplicity.
@@
Through a convolutional structure, a chain of multiple identical protographs can be organized to derive a protograph-based SC-LDPC code with a coupling width of $\omega$. To this end, after edge spreading at each spatial position, $\mathbf{B}$ is split into $\omega+1$ sub-matrices $\mathbf{B}_{i}$$i\in[0,\omega]$, i.e., $\mathbf{B}=\sum_{i=0}^{\omega}\mathbf{B}_{i}$, each edge in the original $\mathbf{B}$ is assigned to exactly one of the sub-matrices. To construct a protomatrix $\mathbf{B}_{\text{SC}}$ of a finite-length SC-LDPC code, $L$ replicas of a series of sub-matrices $\mathbf{B}_{i}$$i\in[0,\omega]$, are coupled to generate $\mathbf{B}_{\text{SC}}$ and the corresponding protograph $\mathcal{G}_{\text{SC}}$, where $L$ is the coupling length. Consequently, a PCM $\mathbf{H}_{\text{SC}}$ can also be derived from a series of sub-PCMs $\mathbf{H}_{i}$$i\in[0,\omega]$ with a lifting size $Z$. Let $\bm{c}_{[0,L-1]}\triangleq[\bm{c}_{0},\bm{c}_{1},\dots,\bm{c}_{L-1}]$ denote a codeword of the SC-LDPC code, and $\bm{c}_{[i]}\triangleq[\bm{u}_{[i]},\bm{r}_{[i]}]$ is the codeword of the $i$-th block (also called batch) in an HARQ process, where $\bm{u}_{[i]}$ is the information vector and $\bm{r}_{[i]}$ is the corresponding parity-check vector. Note that the above $\mathbf{B}_{\text{SC}}$, $\mathcal{G}_{\text{SC}}$, $\mathbf{H}_{\text{SC}}$, and $\bm{c}_{[0,L-1]}$ temporarily ignore the last redundancy tail, which we will explain in detail below.

### v2

@@ Abstract
Next-generation channel coding has stringent demands on throughput, energy consumption, and error rate performance while maintaining key features of 5G New Radio (NR) standard codes such as rate compatibility, which is a significant challenge. Due to excellent capacity-achieving performance, spatially-coupled low-density parity-check (SC-LDPC) codes are considered a promising candidate for next-generation channel coding. In this paper, we propose an SC-LDPC code family called edge-spreading Raptor-like (ESRL) codes. Unlike other SC-LDPC codes that adopt the structure of existing rate-compatible LDPC block codes before coupling, ESRL codes maximize the possible locations of edge placement and focus on constructing an optimal coupled matrix. Moreover, a new graph representation called the unified graph is introduced. This graph offers a global perspective on ESRL codes and identifies the optimal edge reallocation to optimize the spreading strategy. We conduct comprehensive comparisons of ESRL codes and 5G-NR LDPC codes. Simulation results demonstrate that when all decoding parameters and complexity are the same, ESRL codes have obvious advantages in error rate performance and throughput compared to 5G-NR LDPC codes in some specific scenarios (low and high number of iterations), making them a promising solution towards next-generation channel coding.
@@ I Introduction
Next-generation channel coding for enhanced mobile broadband+ (eMBB+) scenarios with high-speed data services targets a peak data rate of $200$ Gbps to $1$ Tbps with an energy efficiency target of around $1$ pJ/bit [1]. This ambitious target represents a 10×10$\times$10 × to 100×100$\times$100 × more stringent throughput requirements compared to 5G New Radio (NR) [2]. As standards continue to evolve, channel coding also needs to meet diverse key performance indicators (KPIs), including higher throughput, enhanced reliability, lower power consumption, reduced computational complexity, rate compatibility, and new capabilities [3], leading to a significant challenge.
@@
As a key technique for eMBB scenarios of 5G-NR [4], low-density parity-check (LDPC) codes [5] have achieved significant success due to capacity-approaching performance [6] and parallel processing capabilities under belief propagation (BP) decoding [7]. For finite-length LDPC codes, though BP offers the advantage of reduced decoding complexity, the BP decoding threshold is sub-optimal compared to the maximum a-posteriori probability (MAP) threshold. Hence, spatially-coupled LDPC (SC-LDPC) codes (also called LDPC convolutional codes) were proposed in [8] and [9], by superimposing LDPC block codes with a convolutional structure. In general, block codes have a simple structure for which decoding is easy to implement in hardware, but they have increased complexity and latency when handling frames with large blocklengths. In contrast, convolutional codes are effective at decoding long frames but suffer from sequential processing of windowed decoding that introduces decoding latency. SC-LDPC codes, offering the best of both worlds, achieve the MAP threshold of the underlying block ensembles using BP decoding on a coupled ensemble [9, 10]. Besides, the compact and regular structure of underlying blocks also simplifies the decoding of long frames, and state-of-the-art SC-LDPC decoder implementations have reported high throughput in the range of hundreds of Gbps [11]. Note that as stated in [12], the information throughput of a decoder depends on clock frequency, achievable parallelism in hardware, code rate, iteration count, and frame length. In general, clock frequency and hardware parallelism are decided by process technology and area constraints. Given the slowdown in frequency scaling in advanced technology nodes, extending the frame length becomes a natural choice for 6G to improve peak throughput [3] and even efficiency. The adjustable coupling length of SC-LDPC codes is therefore a desirable property for 6G eMBB+ to flexibly adjust and extend the frame length without altering the matrix of underlying codes. These ideas have attracted widespread interest for 6G eMBB+ [3, 13].
@@
For 6G eMBB+, rate compatibility and blocklength scalability are also necessary features for channel coding. To address the time-varying nature of wireless channels, rate-compatible channel codes must further support incremental redundancy transmission, thereby maximizing the throughput under hybrid automatic repeat-request (HARQ) schemes [14]. Specifically, a rate-compatible code family is a nested sequence of codes with different rates, where higher rate codes are embedded into lower rate codes [15]. An initial codeword is transmitted at a high rate, and in case of failed decoding, additional partial redundancy is sent until all redundancy of the mother code is used up. Numerous works on the construction of rate-compatible LDPC codes have been published such as [15, 16, 17, 18, 19, 20] mainly using puncturing and extension. The design of such codes can begin with the construction of a relatively small bipartite graph called the protograph [21]. A well-designed protograph applies a copy-and-permute operation (also called lifting) to obtain larger graphs with nodes, thus leading to longer-blocklength LDPC codes [22]. In this context, a protograph-based Raptor-like (PBRL) LDPC code family that meets the above requirements is a very popular one [18]. Sharing many similarities with LDPC codes, the introduction of Luby transform (LT) codes [23] and Raptor codes [24], which can virtually generate any desired rate, form a novel coding paradigm of rateless codes. Subsequently, PBRL codes restrict the protograph to a Raptor-like structure, achieving both rate compatibility and blocklength adjustment simultaneously. Hence, PBRL LDPC codes have been adopted into 5G-NR successfully [4]. The authors of [25, 26] further optimized 5G-NR LDPC codes to enhance the error rate performance for small blocklengths, as they are used for ultra-reliable low-latency communications (URLLC). However, throughput and error floor targets of this mode are very different to eMBB+. In fact, the complexity of BP decoding and hardware increases as the blocklength grows in 5G-NR and for 6G eMBB+. As indicated by the state-of-the-art 5G-NR LDPC decoder implementation [27], the memory overhead required to accommodate the maximum blocklength occupies a significant part of the decoder core area (over $60\%$) and results in complex routing. This limits peak throughput to around $20$ Gbps for 5G-NR LDPC codes, far away from stringent demands beyond 5G. Given the aforementioned advantages of SC-LDPC codes over traditional LDPC block codes, developing a rate-compatible SC-LDPC code family has emerged as a promising solution towards next-generation channel coding for 6G eMBB+.
@@
SC-LDPC codes can be constructed from LDPC block codes using an edge-spreading technique [28]. Several SC-LDPC code constructions tailored to a specific rate were presented in [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], where most of them focus on optimizing edge spreading starting from a given LDPC block code, mitigating the emergence of harmful graph structures. For instance, [29] proposed a heuristic search combined with a check-and-flip process to construct short-cycle-free SC-LDPC codes by generating multiple auxiliary matrices. However, the pseudo-random nature of [29] leads to significant complexity, particularly as the matrix size and coupling width increase. To address this issue, an edge-spreading matrix [31] has been employed in recent works [32, 33, 34, 35, 36, 37, 38] to provide a more compact representation of edge spreading. This enables a simple method to identify harmful structures in SC-LDPC codes, regardless of the increase in coupling width. Naseri et al. [34] imposed constraints on the edge-spreading matrix to reduce dominant trapping sets and avoid error floors[35] developed a strategy to analyze combinations of harmful objects in the matrix, [36] employed gradient descent to reduce the computational complexity of constructing SC-LDPC codes with a large coupling width, and [37] analyzed the relationship between short cycles and edge spreading to design short-cycle-free SC-LDPC codes with minimal coupling width. To further incorporate rate compatibility [40, 41, 42, 43, 44], Si et al. introduced the first rate-compatible SC-LDPC code to achieve the capacity of the binary erasure channel (BEC) [40]. Zhou et al. [41] designed a family of robust rate-compatible punctured SC-LDPC codes by ensuring the recoverability of punctured bits and avoiding harmful graph structures. Protograph-based rate-compatible SC-LDPC codes were proposed in [42, 43, 44] following a simple idea. Namely, a conventional PBRL protograph is first designed, and then the edge-spreading technique is directly applied to construct a coupled code, with both procedures being independent. Notably, [44] achieved performance gains by simply splitting the 5G-NR LDPC matrix as the foundation for SC-LDPC codes.
@@
However, these SC-LDPC works either focus on optimizing code constructions tailored to a specific rate [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] or apply the structure of existing rate-compatible LDPC block codes to generate SC-LDPC codes [42, 43, 44], thus not exploring the full potential of SC-LDPC codes. Consequently, there is a need for a holistic design method to construct SC-LDPC codes that simultaneously meet diverse demands of 6G eMBB+, including rate compatibility, protograph design with coupling, threshold analysis, removal of short cycles, and high throughput potential. In particular, in high-throughput 6G eMBB+ scenarios, the iteration count significantly impacts throughput, where the constructed SC-LDPC codes need to achieve strong error rate performance within just a few iterations. It is also necessary to conduct comprehensive comparisons of the designed SC-LDPC codes and standard LDPC codes, especially when all decoding parameters and complexity are consistent. This would provide a clear assessment of the potential of SC-LDPC codes for next-generation channel coding.
@@
Contributions: This paper proposes an SC-LDPC code family called edge-spreading Raptor-like (ESRL) LDPC codes. Our codes target the 6G eMBB+ mode which requires large and scalable transport block sizes with scalable code rates and support for incremental redundancy HARQ (IR-HARQ), making them well-suited for very high-throughput decoding. Unlike PBRL codes [18], ESRL codes maximize the possible locations of edge placement before coupling and focus on constructing an optimal coupled matrix. Hence, ESRL codes show good performance, while maintaining extensive rate compatibility. The structure, properties, and code profile of ESRL codes are discussed in detail.
Moreover, we propose a novel graph representation called the unified graph. Unlike other edge-spreading optimizations involved with complex processing [34], this graph innovatively offers a global perspective on ESRL codes, identifying the optimal edge reallocation to optimize the spreading strategy. We also introduce modified coupled reciprocal channel approximation (RCA) to analyze the decoding threshold of ESRL codes, demonstrating that ESRL codes enjoy better thresholds across a wide range of rates. Based on these tools, a complete design flow of ESRL codes is presented.
We also propose a high-throughput decoding algorithm tailored to ESRL codes, called semi-layered multi-engine (SLME) decoding. Based on features of ESRL codes, SLME decoding decomposes the original BP decoding into multiple identical sub-decoders that operate in parallel in a more compact manner compared to [13]. This approach significantly enhances the decoding parallelism for ESRL codes.
Finally, we provide a design example of ESRL codes, which targets 6G eMBB+, and conduct comprehensive comparisons with 5G-NR LDPC codes [4] and their variants. As summarized in Fig. 1, ESRL codes preserve several advantages and properties of 5G-NR LDPC codes while further enhancing error rate performance and throughput. When all decoding parameters and complexity are the same, at the rate of $0.52$, ESRL codes outperform 5G-NR LDPC codes by $0.95$ dB at five iterations or enhance peak throughput by $50.1\%$ using comparable hardware complexity [27].
Outline: The remainder of this paper is organized as follows. Section II introduces the symbol definitions and reviews in compact form the preliminaries of LDPC codes. Section III introduces the structure, properties, and code profile of ESRL codes. Section IV outlines the design and optimization procedures for ESRL codes, including the unified graph and modified coupled RCA. In Section V, we present the puncturing scheme together with the integration in the IR-HARQ procedure and SLME decoding for ESRL codes. Section VI provides a design example of ESRL codes and compares it with 5G-NR LDPC codes. Finally, Section VII concludes this paper.
Figure 2: Uncoupled protograph and protomatrix for an ESRL code with an HRC, an IRC, and an SSC. The upper right corner is left empty to ensure rate compatibility.
@@ II Preliminaries
LDPC block codes are specified by a sparse $M\times N$ parity-check matrix (PCM) $\mathbf{H}$, which can be further described by a bipartite Tanner graph [45] with the information length of $K=N-M$. If $\mathbf{H}_{c,v}=1$ for $0\leq c<M$ and $0\leq v<N$, the $c$-th check node (CN) is connected to the $v$-th variable node (VN) on the Tanner graph. We use $\mathbb{V}_{c}$ to represent the set of adjacent VNs for the $c$-th CN and we use $\mathbb{C}_{v}$ to denote the set of neighbors for the $v$-th VN. Let $|\mathbb{V}_{c}|=d_{c}$ and $|\mathbb{C}_{v}|=d_{v}$ refer to the row degree of the $c$-th CN and the column degree of the $v$-th VN, respectively. Due to the quasi-cyclic (QC) property [22], a QC-LDPC code can be described by a more structured $m\times n$ protomatrix $\mathbf{B}$ with the information length of $k=n-m$, where each entry $\mathbf{B}_{c,v}$ can be extended by a $Z\times Z$ cyclic-shifting identity matrix, with $Z$ being the lifting size. To focus on the protomatrix design in this paper, $\mathbf{B}_{c,v}$ denotes only the number of connected edges between the $c$-th CN and the $v$-th VN instead of a specific cyclic shifting value. For 5G-NR LDPC codes [4], the Raptor-like structure of $\mathbf{B}$ is shown in (1), which consists of three components to achieve rate compatibility [18], i.e., a highest-rate code (HRC) $\mathbf{B}_{\text{HRC}}$, an incremental redundancy code (IRC) $\mathbf{B}_{\text{IRC}}$, and an identity matrix $\mathbf{I}$.
$\mathbf{B}=\left[\begin{array}[]{c c}\mathbf{B}_{\text{HRC}}&\mathbf{0}\\ \mathbf{B}_{\text{IRC}}&\mathbf{I}\\ \end{array}\right].$ (1)
Similarly, $\mathbf{B}$ can be transformed into a protograph $\mathcal{G}$ [21], defined as $\mathcal{G}\triangleq\{\mathbb{V}\cup\mathbb{C},\mathbb{E}\}$, where $\mathbb{E}$ is the set of all edges in $\mathcal{G}$. Let $e_{c}^{v}$ denote the directed edge from the $c$-th CN to the $v$-th VN, and $e_{v}^{c}$ is the same edge with the opposite direction. Note that all graph concepts are based on protographs throughout this paper and we omit the term ‘QC’ for simplicity.
@@
Through a convolutional structure, a chain of multiple identical protographs can be organized to derive a protograph-based SC-LDPC code with a coupling width of $\omega$. To this end, after edge spreading at each spatial position, $\mathbf{B}$ is split into $\omega+1$ sub-matrices $\mathbf{B}_{i}$$i\in[0,\omega]$, i.e., $\mathbf{B}=\sum_{i=0}^{\omega}\mathbf{B}_{i}$, each edge in the original $\mathbf{B}$ is assigned to exactly one of the sub-matrices. To construct a protomatrix $\mathbf{B}_{\text{SC}}$ of a finite-length SC-LDPC code, $L$ replicas of a series of sub-matrices $\mathbf{B}_{i}$$i\in[0,\omega]$, are coupled to generate $\mathbf{B}_{\text{SC}}$ and the corresponding protograph $\mathcal{G}_{\text{SC}}$, where $L$ is the coupling length. Consequently, a PCM $\mathbf{H}_{\text{SC}}$ can also be derived from a series of sub-PCMs $\mathbf{H}_{i}$$i\in[0,\omega]$ with a lifting size $Z$. Let $\bm{c}_{[0,L-1]}\triangleq[\bm{c}_{0},\bm{c}_{1},\dots,\bm{c}_{L-1}]$ denote a codeword of the SC-LDPC code, and $\bm{c}_{[i]}\triangleq[\bm{u}_{[i]},\bm{r}_{[i]}]$ is a partial codeword of the $i$-th block (also called batch) in an IR-HARQ process, where $\bm{u}_{[i]}$ is the information vector and $\bm{r}_{[i]}$ is the corresponding parity-check vector. Note that the above $\mathbf{B}_{\text{SC}}$, $\mathcal{G}_{\text{SC}}$, $\mathbf{H}_{\text{SC}}$, and $\bm{c}_{[0,L-1]}$ temporarily ignore the last redundancy tail, which we will explain in detail below.